# Gappy Data Reconstruction using Unsupervised Learning for Digital Twin

### 13 Dec 2023 v1

@@ 1 Introduction
Methods to find a suitable generalized coordinate can be categorized into two approaches: intrusive and non-intrusive. In intrusive methods, the governing equations are solved with a substituted solution approximation. Projection-based reduced-order models, such as the linear subspace reduced-order model [7, 16, 12, 15, 17, 10], nonlinear manifold reduced-order model [36, 30, 29, 19], component-wise reduced-order model [38, 39] and space-time reduced-order model [14, 13, 32, 47], are examples of intrusive methods. These techniques are widely employed to expedite physical simulations.
In contrast, non-intrusive methods do not involve solving governing equations. Dynamic identification (DI) methods presented in [23, 26, 3] identify dynamics in latent space using regression method in a least–squares sense. The Gappy proper orthogonal decomposition (POD) method developed by [22] approximates a solution with a linear subspace spanned by the POD basis matrix. The gappy POD method finds generalized coordinates, which are POD basis coefficients that minimize the error between the predicted and gappy data. This method was successfully applied to unsteady flow reconstruction [48]. Linear subspace-based approaches can represent data with fast decaying Kolmogorov N-width, such as elliptic and parabolic equations. However, such methods are not effective in representing data with slowly decaying Kolmogorov N-width, such as linear transport problems [43] and wave equation [24]. The dynamic mode decomposition (DMD) method is a data-driven and non-intrusive method developed by Peter Schmid [45, 34] that approximates local dynamic systems from the dynamic modes found using eigenvectors and eigenvalues of the reduced system obtained through singular value decomposition (SVD) of the dynamic system.
@@
There is also an auto-encoder-based approach to reconstruct data. When gappy data is given as an input, a model outputs reconstructed data directly. In [25], the authors presented a conditional variational auto-encoder (CVAE)-based method by conditioning on measurements, simultaneously predicting the solution and uncertainty. Their model reduces the input data to latent space variables, which describe probability distributions. The probability distribution in the late space is found with the measurement data. Then, solutions can be produced from observations and latent space variables. In [42], a deep neural network is used to predict the reduced state of full-order data from sensor measurements directly. Then, the full-order data is reconstructed by multiplying POD basis matrix by the reduced state. In [25] and [42], measurement data is an input argument of the neural networks and is used during the forward pass during inference time. However, our proposed method does not take measurement data as the input of the neural network model. In our method, we solve the minimization problem between the measurement data and the model prediction to find generalized coordinates in a latent space.
We developed a novel gappy data reconstruction method called Gappy auto-encoder (AE) to represent physics field data for digital twin purposes. The proposed method learns a nonlinear mapping between the latent and original data spaces using an auto-encoder. Then, we find generalized coordinates in the latent space by minimizing the error in $l_{2}$ norm between gappy measured data and the decoder output. It is categorized into a data-driven and non-intrusive method. Compared to the Gappy POD method, our Gappy AE method represents data using a nonlinear manifold, overcoming the limitations of linear subspace-based methods. The major contributions of our method are summarized as follows:
@@ 2.1 Gappy POD
(2.3) $\boldsymbol{\Phi}=\underset{}{\text{argmin}}\|\boldsymbol{X}-\tilde{\boldsymbol{\Phi}}\tilde{\boldsymbol{\Phi}}^{T}\boldsymbol{X}\|_{F}^{2},$
where $\|\boldsymbol{A}\|_{F}$ denotes the Frobenius norm of a matrix $\boldsymbol{A}$. For more detailed information on the POD method, please refer [27, 33].
We can set a minimization problem between the measurement and solution data approximation given the sparse measurement data. By solving the minimization problem, we find the generalized coordinates, $\hat{\boldsymbol{x}}$. The Gappy POD first approximates $\boldsymbol{x}\in\mathbb{R}^{N_{s}}$ as
@@
where $\boldsymbol{\Phi}\in\mathbb{R}^{N_{s}\times{n_{s}}}$ is the POD basis matrix from the solution snapshot matrix, assuming $\boldsymbol{x}_{ref}=\mathbf{0}$. Solving the following minimization problem
(2.5) $\lVert\boldsymbol{Z}^{T}(\boldsymbol{x}-\boldsymbol{\Phi}\hat{\boldsymbol{x}})\rVert_{2}^{2}$
@@
(2.6) $\boldsymbol{x}\approx\boldsymbol{\Phi}(\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{\dagger}\boldsymbol{Z}^{T}\boldsymbol{x},$
where ${}^{\dagger}$ denotes the Moore–Penrose inverse of a matrix and $\boldsymbol{Z}^{T}\boldsymbol{x}$ is replaced with the sparse measurement data. The graphical summary of Gappy POD process is shown in Fig. 1.
@@ 2.2.2 Sparse Shallow Auto-Encoder
Auto-Encoder is a feed-forward neural network model that learns identity mapping. The auto-encoder consists of an encoder, $\boldsymbol{h}:\mathbb{R}^{N_{s}}\rightarrow\mathbb{R}^{{n_{s}}}$, and a decoder, $\boldsymbol{g}:\mathbb{R}^{{n_{s}}}\rightarrow\mathbb{R}^{N_{s}}$. The nonlinear activation functions in the auto-encoder make them nonlinear. The encoder maps high dimensional data, $\boldsymbol{x}\in\mathbb{R}^{N_{s}}$ to low dimensional data, $\hat{\boldsymbol{x}}\in\mathbb{R}^{{n_{s}}}$ and the decoder maps back the low dimensional data to the high dimensional data. The decoder is used for nonlinear mapping mentioned in Section 2.2.1.
The general structure of the encoder and decoder can vary, but for our study, we use a sparse shallow network as outlined in Section 3.2 of [30]. The sparsity in the network enhances efficiency by eliminating connections in the layers. However, when constructing neural networks for physical data with higher dimensions, caution must be exercised to ensure that the spatial data organization aligns with the sparse connection. In other words, spatially near nodes share part of the connections, whereas far-away nodes in the space domain do not share the connections even if two nodes are neighbors. In this work, the encoder and decoder have three layers. The first layers are fully connected with nonlinear activation functions, while the final layer lacks nonlinear activation functions. The masking operator is then applied to induce sparsity in the decoder. For more information on the neural network architecture, please refer to [30].
@@ 2.2.3 Gappy Procedures
(2.8) $n_{i}=\begin{cases}1,&\text{if $u_{i}$ is known}\\ 0,&\text{if $u_{i}$ is missing}\end{cases}$
where $\mathbf{u}\in\mathbb{R}^{N_{s}}$ is a vector that represents the solution at a given time and parameter. Then, a diagonal matrix, $\mathbf{N}\in\mathbb{R}^{N_{s}\times N_{s}}$ is constructed with elements of $\mathbf{n}$ vector on the diagonal and removing zero columns gives us $\mathbf{Z}\in\mathbb{R}^{N_{s}\times n_{r}}$, where $n_{r}$ is the number of known values. Now, we define an error, $\epsilon\in\mathbb{R}$ as
(2.9) $\epsilon=\lVert\boldsymbol{Z}^{T}(\mathbf{u}-\tilde{\mathbf{u}})\rVert_{2}^{2},$
where $\boldsymbol{Z}^{T}\mathbf{u}$ is a measurement data and $\tilde{\mathbf{u}}$ is prediction given by
(2.10) $\tilde{\mathbf{u}}=\mathbf{u}_{ref}+\boldsymbol{g}(\mathbf{b}),$
where $\boldsymbol{g}:\mathbb{R}^{{n_{s}}}\rightarrow\mathbb{R}^{N_{s}}$ with ${n_{s}}\ll N_{s}$ denotes a nonlinear function that maps a latent space dimension ${n_{s}}$ to the full order model space dimension $N_{s}$; $\mathbf{b}\in\mathbb{R}^{{n_{s}}}$ is a generalized coordinate to be found.
Minimizing $\epsilon$ gives us a generalized coordinate $\mathbf{b}$ through the Gauss–Newton method. At the $i$th iteration, first, we linearize $\boldsymbol{Z}^{T}(\mathbf{u}-\tilde{\mathbf{u}})$ at $\mathbf{b}_{i}$ as
(2.11) $\boldsymbol{Z}^{T}(\mathbf{u}-\mathbf{u}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})-\mathbf{J}\Delta\mathbf{b}_{i}),$
where $\mathbf{J}=\frac{\partial\tilde{\mathbf{u}}}{\partial\mathbf{b}}$ and $\Delta\mathbf{b}_{i}=\mathbf{b}_{i+1}-\mathbf{b}_{i}$. Then, the linearized error $\epsilon_{lin}$ becomes
(2.12) $\epsilon_{lin}=(\mathbf{u}-\mathbf{u}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})-\mathbf{J}\Delta\mathbf{b}_{i})^{T}\boldsymbol{Z}\boldsymbol{Z}^{T}(\mathbf{u}-\mathbf{u}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})-\mathbf{J}\Delta\mathbf{b}_{i}).$
@@
(2.13) $\frac{\partial\epsilon_{lin}}{\partial\mathbf{b}}$ $=2\mathbf{J}^{T}\boldsymbol{Z}\boldsymbol{Z}^{T}(\mathbf{u}-\mathbf{u}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})-\mathbf{J}\Delta\mathbf{b}_{i})$
@@
(2.15) $\Delta\mathbf{b}_{i}$ $=(\mathbf{J}^{T}\boldsymbol{Z}\boldsymbol{Z}^{T}\mathbf{J})^{-1}\mathbf{J}^{T}\boldsymbol{Z}\boldsymbol{Z}^{T}(\mathbf{u}-\mathbf{u}_{ref}-\boldsymbol{g}(\mathbf{b}_{i}))$
(2.16) $=(\boldsymbol{Z}^{T}\mathbf{J})^{\dagger}\boldsymbol{Z}^{T}(\mathbf{u}-\mathbf{u}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})).$
@@
(2.17) $\mathbf{b}_{i+1}=\mathbf{b}_{i}+(\boldsymbol{Z}^{T}\mathbf{J})^{\dagger}\boldsymbol{Z}^{T}(\mathbf{u}-\mathbf{u}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})).$
The above steps are iterated until the error is minimized. Let us define $\mathbf{b}_{f}$ as the generalized coordinate at the final iteration. Then $\tilde{\mathbf{u}}$ is computed using $\tilde{\mathbf{u}}=\mathbf{u}_{ref}+\boldsymbol{g}(\mathbf{b}_{f})$ to complete the reconstruction.

### 31 Mar 2024 v2

# Gappy AE: A Nonlinear Approach for Gappy Data Reconstruction using Auto-Encoder

@@ 1 Introduction
Methods to find a suitable generalized coordinate can be categorized into two approaches: intrusive and non-intrusive. In intrusive methods, the governing equations are solved with a substituted solution approximation. Projection-based reduced-order models, such as the linear subspace reduced-order model [7, 16, 12, 15, 18, 10], nonlinear manifold reduced-order model [37, 31, 30, 20], component-wise reduced-order model [39, 40] and space-time reduced-order model [14, 13, 33, 49], are examples of intrusive methods. These techniques are widely employed to expedite physical simulations. Recent research in reduced order models (ROMs) focuses on enhancing the representability of autoencoders [46] and the computational efficiency of autoencoder training [17, 20]. In [17], the authors reduced computational costs by training autoencoders on subsampled points only. During the online phase, the decoder outputs data corresponding to subsampled points, and then a POD basis matrix is used to reconstruct the data. The authors in [20] present the domain decomposition (DD) method for nonlinear-manifold reduced order models (NM-ROMs). In NM-ROMs, nonlinear manifolds are found by training sparse and shallow autoencoders, which can be costly for large-scale problems. To alleviate the training cost of the sparse and shallow autoencoders, the authors decomposed the domain into subdomains, resulting in the reduction of the number of learnable parameters of autoencoders for each subdomain and the size of the training data. In [46], the authors introduced convolutional decoders and compressed decoders. The convolutional decoder learns a latent representation of the dynamical system. There is no constraint on the structure of the convolutional decoder, giving us the opportunity to test various deep learning models. The compressed decoder learns mapping from latent space to subsampled data. The latent space coordinates are found by minimizing subsampled residuals, and then the latent space coordinates are taken as input for the convolutional decoder to predict the solution. In this case, subsampled residuals are only used to find latent space coordinates, and the convolutional decoder computes full-order outputs.
In contrast, non-intrusive methods do not involve solving governing equations. Dynamic identification (DI) methods presented in [24, 27, 3] identify dynamics in latent space using regression method in a least–squares sense. The Gappy proper orthogonal decomposition (POD) method developed by [23] approximates a solution with a linear subspace spanned by the POD basis matrix. The gappy POD method finds generalized coordinates, which are POD basis coefficients that minimize the error between the predicted and gappy data. This method was successfully applied to unsteady flow reconstruction [50]. Linear subspace-based approaches can represent data with fast decaying Kolmogorov N-width, such as elliptic and parabolic equations. However, such methods are not effective in representing data with slowly decaying Kolmogorov N-width, such as linear transport problems [44] and wave equation [25]. For digital twins of problems with Kolmogorov N-width decaying slowly, e.g., construction of earthquake digital twin model for source characterization, we need more general method. Since seismic waves are solutions to the wave equation, conventional Gappy POD does not perform well in terms of data reconstruction. The dynamic mode decomposition (DMD) method is a data-driven and non-intrusive method developed by Peter Schmid [47, 35] that approximates local dynamic systems from the dynamic modes found using eigenvectors and eigenvalues of the reduced system obtained through singular value decomposition (SVD) of the dynamic system.
@@
There is also an auto-encoder-based approach to reconstruct data. When gappy data is given as an input, a model outputs reconstructed data directly. In [26], the authors presented a conditional variational auto-encoder (CVAE)-based method by conditioning on measurements, simultaneously predicting the solution and uncertainty. Their model reduces the input data to latent space variables, which describe probability distributions. The probability distribution in the latent space is found with the measurement data. Then, solutions can be produced from observations and latent space variables. In [43], a deep neural network is used to predict the reduced state of full-order data from sensor measurements directly. Then, the full-order data is reconstructed by multiplying POD basis matrix by the reduced state. In [26] and [43], measurement data is an input argument of the neural networks and is used during the forward pass during inference time. However, our proposed method does not take measurement data as the input of the neural network model. In our method, we solve the minimization problem between the measurement data and the model prediction to find generalized coordinates in a latent space.
We developed a novel gappy data reconstruction method called Gappy auto-encoder (AE) to represent physics field data for digital twin purposes. The proposed method learns a nonlinear mapping between the latent and original data spaces using an auto-encoder. Then, we find generalized coordinates in the latent space by minimizing the error in $l_{2}$ norm between gappy measured data and the decoder output. It is categorized into a data-driven and non-intrusive method. Compared to the Gappy POD method, our Gappy AE method represents data using a nonlinear manifold, overcoming the limitations of linear subspace-based methods. Replacing a linear subspace basis matrix with an auto-encoder requires special care, which is not trivial, such as determining sparse sampling locations, having sparse and shallow decoder structure, and creating the subnet that only contains active nodes and weights as detailed in Section 3 and 4 of [31]. The purpose of these non-trivial things is to compute the Jacobian of the decoder quickly which is used in Gauss-Newton iterations for finding generalized coordinates in a latent space. The major contributions of our method are summarized as follows:
@@
Nomenclature
  • $N_{s}$
    Full order model space dimension
  • ${n_{s}}$
    Reduced space dimension
  • $n_{r}$
    The number of samples
  • $n_{\mu}$
    The number of parameters
  • ${N_{t}}$
    The number of time steps
  • $\boldsymbol{x}\in\mathbb{R}^{N_{s}}$
    Solution
  • $\boldsymbol{x}_{ref}\in\mathbb{R}^{N_{s}}$
    Reference solution
  • $\hat{\boldsymbol{x}}\in\mathbb{R}^{{n_{s}}}$
    Generalized coordinates, or reduced solution
  • $\boldsymbol{U}\in\mathbb{R}^{N_{s}\times n_{\mu}({N_{t}}+1)}$
    Left singular matrix
  • $\boldsymbol{V}\in\mathbb{R}^{n_{\mu}({N_{t}}+1)\times n_{\mu}({N_{t}}+1)}$
    Right singular matrix
  • $\boldsymbol{\Sigma}\in\mathbb{R}^{n_{\mu}({N_{t}}+1)\times n_{\mu}({N_{t}}+1)}$
    Singular values matrix
  • $\boldsymbol{\Phi}\in\mathbb{R}^{N_{s}\times{n_{s}}}$
    Basis matrix with ${n_{s}}\ll N_{s}$
  • $\boldsymbol{g}:\mathbb{R}^{{n_{s}}}\rightarrow\mathbb{R}^{N_{s}}$
    Nonlinear function with ${n_{s}}\ll N_{s}$
  • $\lVert\boldsymbol{T}\rVert_{F}$
    Frobenius norm of a matrix $\boldsymbol{T}$
  • $\boldsymbol{T}^{\dagger}$
    Moore–Penrose inverse of a matrix $\boldsymbol{T}$
  • $\mathbf{J}(\hat{\boldsymbol{x}})$
    Jacobian of $\boldsymbol{g}(\hat{\boldsymbol{x}})$
  • $\boldsymbol{Z}^{T}\in\mathbb{R}^{n_{r}\times N_{s}}$
    Sampling matrix with $n_{r}\geq{n_{s}}$
  • $\mathbf{r}\in\mathbb{R}^{N_{s}}$
    Residual
  • $\hat{\mathbf{r}}\in\mathbb{R}^{n_{r}}$
    Reduced residual
  • $\boldsymbol{\Phi}_{r}\in\mathbb{R}^{N_{s}\times n_{r}}$
    Residual POD basis matrix
@@ 2.1 Gappy POD
(2.3) $\boldsymbol{\Phi}=\underset{}{\text{argmin}}\|\boldsymbol{X}-\tilde{\boldsymbol{\Phi}}\tilde{\boldsymbol{\Phi}}^{T}\boldsymbol{X}\|_{F}^{2}.$
For more detailed information on the POD method, please refer to [28, 34].
For brevity and readability, $n$ and $\boldsymbol{\mu}_{k}$ are omitted in the notation $\boldsymbol{x}_{n}^{\boldsymbol{\mu}_{k}}$. We can set a minimization problem between the measurement and solution data approximation given the sparse measurement data. By solving the minimization problem, we find the generalized coordinates, $\hat{\boldsymbol{x}}$. The Gappy POD first approximates $\boldsymbol{x}\in\mathbb{R}^{N_{s}}$ as
@@
assuming $\boldsymbol{x}_{ref}=\mathbf{0}$. Solving the following minimization problem
(2.5) $\hat{\boldsymbol{x}}=\underset{\hat{\boldsymbol{v}}\in\mathbb{R}^{{n_{s}}}}{\text{argmin}}\lVert\boldsymbol{Z}^{T}(\boldsymbol{x}-\boldsymbol{\Phi}\hat{\boldsymbol{v}})\rVert_{2}^{2}$
@@
(2.6) $\hat{\boldsymbol{x}}=(\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{\dagger}\boldsymbol{Z}^{T}\boldsymbol{x},$
and then we have
(2.7) $\boldsymbol{x}\approx\boldsymbol{\Phi}\hat{\boldsymbol{x}}=\boldsymbol{\Phi}(\boldsymbol{Z}^{T}\boldsymbol{\Phi})^{\dagger}\boldsymbol{Z}^{T}\boldsymbol{x},$
where $\boldsymbol{Z}^{T}\boldsymbol{x}$ is replaced with the sparse measurement data. The graphical summary of Gappy POD process is shown in Fig. 1.
@@ 2.2.2 Sparse Shallow Auto-Encoder
Auto-Encoder is a feed-forward neural network model that learns identity mapping. The auto-encoder consists of an encoder, $\boldsymbol{h}:\mathbb{R}^{N_{s}}\rightarrow\mathbb{R}^{{n_{s}}}$, and a decoder, $\boldsymbol{g}:\mathbb{R}^{{n_{s}}}\rightarrow\mathbb{R}^{N_{s}}$. The nonlinear activation functions in the auto-encoder make them nonlinear. The encoder maps high dimensional data, $\boldsymbol{x}\in\mathbb{R}^{N_{s}}$ to low dimensional data, $\hat{\boldsymbol{x}}\in\mathbb{R}^{{n_{s}}}$ and the decoder maps back the low dimensional data to the high dimensional data. The decoder is used for nonlinear mapping, $\boldsymbol{g}(\hat{\boldsymbol{x}})$, mentioned in Section 2.2.1.
The general structure of the encoder and decoder can vary, but for our study, we use a sparse shallow network as outlined in Section 3.2 of [31]. The sparsity in the network enhances efficiency by eliminating connections in the layers. However, when constructing neural networks for physical data with higher dimensions, caution must be exercised to ensure that the spatial data organization aligns with the sparse connection. In other words, spatially near nodes share part of the connections, whereas far-away nodes in the space domain do not share the connections even if two nodes are neighbors. In this work, the encoder and decoder have three layers. The encoder layers are fully connected and a nonlinear activation function is applied in the hidden layer, while the output layer lacks a nonlinear activation function. For the decoder architecture, the nodes between the input and hidden layer are fully connected with a nonlinear activation function, however, the nodes between the hidden and output layer are sparsely connected without an activation function. We have introduced two hyper-parameters, $b$, and $\delta b$, to design a sparsity structure. $b$ denotes the size of a block to compute for one output node and $\delta b$ defines the amount of shift that the block moves. In our work, a swish activation function is used. For more information on the neural network architecture, please refer to [31].
@@ 2.2.3 Gappy Procedures
(2.9) $n_{j}=\begin{cases}1,&\text{if $x_{j}$ is known}\\ 0,&\text{if $x_{j}$ is missing}\end{cases}$
where $\boldsymbol{x}\in\mathbb{R}^{N_{s}}$ is a vector that represents the solution at a given time and parameter and $x_{j}$ denotes $j$-th component of $\boldsymbol{x}$. Then, a diagonal matrix, $\mathbf{N}\in\mathbb{R}^{N_{s}\times N_{s}}$ is constructed with elements of $\mathbf{n}$ vector on the diagonal and removing zero columns gives us $\mathbf{Z}\in\mathbb{R}^{N_{s}\times n_{r}}$, where $n_{r}$ is the number of known values. Now, we define an error, $\epsilon\in\mathbb{R}$ as
(2.10) $\epsilon=\lVert\boldsymbol{Z}^{T}(\boldsymbol{x}-\tilde{\boldsymbol{x}})\rVert_{2}^{2},$
where $\boldsymbol{Z}^{T}\boldsymbol{x}$ is a measurement data and $\tilde{\boldsymbol{x}}$ is prediction given by
(2.11) $\tilde{\boldsymbol{x}}=\boldsymbol{x}_{ref}+\boldsymbol{g}(\mathbf{b}),$
where $\mathbf{b}\in\mathbb{R}^{{n_{s}}}$ is a generalized coordinate to be found.
Minimizing $\epsilon$ gives us a generalized coordinate $\mathbf{b}$ through the Gauss–Newton method. At the $i$-th iteration, first, we linearize $\boldsymbol{Z}^{T}(\boldsymbol{x}-\tilde{\boldsymbol{x}})$ at $\mathbf{b}_{i}$ as
(2.12) $\boldsymbol{Z}^{T}(\boldsymbol{x}-\boldsymbol{x}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})-\mathbf{J}\Delta\mathbf{b}_{i}),$
where $\mathbf{J}=\frac{\partial\tilde{\boldsymbol{x}}}{\partial\mathbf{b}}$ and $\Delta\mathbf{b}_{i}=\mathbf{b}_{i+1}-\mathbf{b}_{i}$. Then, the linearized error $\epsilon_{lin}$ becomes
(2.13) $\epsilon_{lin}=(\boldsymbol{x}-\boldsymbol{x}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})-\mathbf{J}\Delta\mathbf{b}_{i})^{T}\boldsymbol{Z}\boldsymbol{Z}^{T}(\boldsymbol{x}-\boldsymbol{x}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})-\mathbf{J}\Delta\mathbf{b}_{i}).$
@@
(2.14) $\frac{\partial\epsilon_{lin}}{\partial\mathbf{b}}$ $=2\mathbf{J}^{T}\boldsymbol{Z}\boldsymbol{Z}^{T}(\boldsymbol{x}-\boldsymbol{x}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})-\mathbf{J}\Delta\mathbf{b}_{i})$
@@
(2.16) $\Delta\mathbf{b}_{i}$ $=(\mathbf{J}^{T}\boldsymbol{Z}\boldsymbol{Z}^{T}\mathbf{J})^{-1}\mathbf{J}^{T}\boldsymbol{Z}\boldsymbol{Z}^{T}(\boldsymbol{x}-\boldsymbol{x}_{ref}-\boldsymbol{g}(\mathbf{b}_{i}))$
(2.17) $=(\boldsymbol{Z}^{T}\mathbf{J})^{\dagger}\boldsymbol{Z}^{T}(\boldsymbol{x}-\boldsymbol{x}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})).$
@@
(2.18) $\mathbf{b}_{i+1}=\mathbf{b}_{i}+(\boldsymbol{Z}^{T}\mathbf{J})^{\dagger}\boldsymbol{Z}^{T}(\boldsymbol{x}-\boldsymbol{x}_{ref}-\boldsymbol{g}(\mathbf{b}_{i})).$
The above steps are iterated until the error is minimized. Let us write $\hat{\boldsymbol{x}}$ as the generalized coordinate at the final iteration. Then $\tilde{\boldsymbol{x}}$ is computed using $\tilde{\boldsymbol{x}}=\boldsymbol{x}_{ref}+\boldsymbol{g}(\hat{\boldsymbol{x}})$ to complete the reconstruction.